OpenStudy (anonymous):

Which of the following is not in the range of 7(x - 2)2?

OpenStudy (anonymous):

A. -3 B. 12 C. 6 D. 300

OpenStudy (anonymous):

-3 maybe?

OpenStudy (anonymous):

\[7(x-2)^2=7(x-2)(x-2)\]

OpenStudy (anonymous):

find the min and that is your answer

OpenStudy (anonymous):

well yeah that's the answer and the reason for that if because when we expand like I did you can foil and then we get \[7(x^2-4x+4)=7x^2-28x+4\]

OpenStudy (anonymous):

makes sense ! :))

OpenStudy (anonymous):

so if your min is zero and -3 lies below it it cant be part of it.